OpenStudy (anonymous):

You wish to accumulate $100,000 through monthly payments of $100. If you can earn interest at an annual rate of 4% compounded monthly, how long (to the nearest year) will it take to accomplish your goal? I just need the equation, can anyone help me?

ganeshie8 (ganeshie8):

u may use annuity formula

ganeshie8 (ganeshie8):

\(\large 100000 = 100\bullet \frac{(1+i)^{nt}-1}{i}\)

ganeshie8 (ganeshie8):

plugin the known values and solve \(t\)

OpenStudy (anonymous):

first i have to divide 100000/1000 right?

OpenStudy (anonymous):

sorry 100000/100

ganeshie8 (ganeshie8):

yes, did u find \(i\) ?

OpenStudy (anonymous):

yeah it is r/m or 0.04/12

ganeshie8 (ganeshie8):

yes! keep going :)

OpenStudy (anonymous):

I'm sorry, I love math, but sometimes I get confuse about formulas D:, then do i have to multiply 1,000 by .04/12?

ganeshie8 (ganeshie8):

correct !

ganeshie8 (ganeshie8):

you will have to use calculator...

OpenStudy (anonymous):

yeah, but it is the same if first i multiply 1000 by 12 and then divide by .04? or I have to get first the answer of .04/12 ?

ganeshie8 (ganeshie8):

\(\large 100000 = 100\bullet \frac{(1+i)^{nt}-1}{i}\) divide 100 both sides \(\large \frac{100000}{100} = \frac{100\bullet \frac{(1+i)^{nt}-1}{i}}{100}\) \(\large 1000 = \frac{(1+i)^{nt}-1}{i}\) multiply \(i\) both sides : \(\large 1000\times i = (1+i)^{nt}-1\) add 1 both sides : \(\large 1000\times i + 1 = (1+i)^{nt}\)

ganeshie8 (ganeshie8):

now plugin i = 0.04/12

ganeshie8 (ganeshie8):

\(\large 1000\times \frac{0.04}{12} + 1 = (1+\frac{0.04}{12})^{12t}\)

ganeshie8 (ganeshie8):

fine so far eh ?

OpenStudy (anonymous):

Yeah!!! Thanks to be patient, sometimes I just get so confuse and I don't know if there's something wrong with my calculator D:

ganeshie8 (ganeshie8):

your calculation is perfectly fine :)

ganeshie8 (ganeshie8):

wolfram says \( t \approx 36.7\) that means he has to wait \(\approx 37\) years

ganeshie8 (ganeshie8):

see if that looks okay...

OpenStudy (anonymous):

Thank you!!!!!!!!!! You save my life!!

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

Well right now i know how to do the rest of my homework! Thank you so much! (: